# The Merging Galaxy Cluster Environment Affects the Morphology of Radio-AGN

### v1

@@ 2.2 LoTSS Data
To examine the radio characteristics of the cluster members, we used data from the LOFAR Two-metre Sky Survey (LoTSS: Shimwell et al. 2017). LoTSS provides a resolution of 6′′ with a central frequency of 144 MHz and a median RMS sensitivity in public data of 83 Jy/beam. We made use of the current public data release of LoTSS which is Data Release 2 (Shimwell et al., 2022), and in addition we made use of data processed since that release, which will be made publicly available in LoTSS Data Release 3.
Out of the 33 clusters from Bilton et al. (2020), 13 were covered by the LoTSS survey, 9 non-merging and 4 merging. Of these 13 clusters, one cluster, Abell 119, was only half covered by the Sloan Digital Sky Survey and was therefore discarded from the analysis. As a result, our final sample consisted of 12 clusters fully covered by LoTSS and SDSS (see Table 1). For each of these clusters we made a mosaiced image which combined all of the available LoTSS pointings covering that cluster, maximizing the sensitivity of the images used for this study.
@@ 2.3 Completeness and Redshift Distribution
It is known that SDSS completeness drops to 65% in denser fields such as a galaxy cluster (Yoon et al., 2008) due to there being more targets than available fibers. To investigate the completeness of our sample, we compare the average mass of the galaxy clusters and number of cluster members across both dynamical states. When examining the Bilton et al. (2020) SDSS catalogs of the LoTSS covered clusters, we find an average 62 members per merging cluster and 82 per non-merging cluster. When we match these averages with average cluster halo mass, we find mergers have 5.4 x M⊙ and non-mergers 6.3 x M⊙. Cluster mass, M200, is determined in Bilton et al. (2020) from the density profile , where is specifically 200 times the critical density of the flat Universe. If we compare the number of cluster members and halo masses of the clusters, we find merging clusters relative to non-mergers have about 75% of the members and 80% of the mass. The similarity of these values is suggestive that there’s not a significant selection bias.
In addition to survey completeness, we investigate the redshift distribution within each merging and non-merging populations to ensure that they are well-matched. Given this work is rooted in morphology, there exists a potential bias of a morphology class being mistaken for another due to resolution issues coming from redshift differences. For the non-mergers, we find the mean , median , and interquartile range to be 0.06, 0.07, and 0.05. For mergers, we find the mean , median , and interquartile range to be 0.06, 0.06, and 0.05. The redshifts are well matched between the cluster dynamical state populations and we do not expect redshift to bias morphological classifications within each population.
@@ 3 Methods
In order to determine whether or not the merging cluster environment affects the radio morphology of AGN and cluster galaxies, the radio sources need to be classified according to their morphology. To do this, we created images (§3.1), devised a classification system (§3.2), visually inspected the radio sources (§3.3), and determined which sources were radio-AGNs (§3.5).
@@ 3.1 Cutouts
Since there were a large number of sources and respective cutouts to be made, there was a necessity to be able to create bulk cutouts for the classification process. Thus we calculated and used a “global” cluster noise (henceforth referred to as RMS or ) for generating contours for the radio sources. To determine the global RMS for each LoTSS image, we made use of SAOImageDS9 (Joye & Mandel 2003). For each mosaiced cluster image, we selected an emission-free region and drew a circle 4 arcminutes in diameter within this region. From this circle, we extracted the RMS in mJy beam-1 from the statistics calculated automatically in DS9. For each cluster, we used this global cluster RMS to produce the contours for the cluster member cutouts (reported in Table 1).
The method of calculating the RMS and contour levels displayed for the cutouts can influence the appearance and thus the morphology classification of the source. Given our dataset, which comprises over 1000 sources, we were confronted with a trade-off between efficiency and sensitivity. We investigate whether the morphology classification of a source is RMS (and corresponding contour levels) dependent by selecting a handful of sources (14) showing faint emission potentially related to the source but not captured with contours and testing the effect of using different RMS values. For these sources, we created new cutouts using varying RMS and contour levels. We compared the original, global RMS and contour level for each image to 1) the local RMS (determined from a ′′ circle of an emission free region in the cutout) with the original 4 2n base for contours, 2) the original global (cluster) RMS but with a 3, 4, 5, 8 contour levels, and 3) the local RMS with 3, 4, 5, 8 contour levels. We find no significant deviation from the original classification results when comparing the new RMS and contour levels. As a result, we are confident that our choice of using a global RMS is robust.
To aid with the classification process, we displayed contours in the cutouts that were log-based starting with a 4 base determined by the cluster RMS, as described above. Thus, the contours begin with 4 and increase by factors of 2n where n = 1, 2, 3, etc. To optimize visualization for many hundreds of images, we employed a square root scaling technique in aplpy and scaled the images as a percentage of the maximum pixel value (pmax=100). To see all of these parameters in use see Figure 1.
@@ 3.2 Radio Morphology Classifications
After the cutouts were made, one team member visually categorized the sources as having a radio detection or non-detection. A source was defined as having a radio detection if it had 8 detection at the cluster member coordinate from the Bilton et al. (2020) catalog, which manifested visually as two contours (at a 4 base). We chose 8 as the threshold for a detection for the following reasoning. Because we used a global, cluster RMS to create the contours for the cluster members, there are a few cases where 4 contour encapsulates the noise instead of the morphology of the source, making the radio morphology of the source difficult to classify. Since we were classifying the sources as detection or non-detection visually (to be consistent with the visual radio morphology classification), we increased the base of detection to the next displayed contour which in this case is 8 (because we define the contours as 4 2n where n = 0, 1, 2, etc.) in order to be confident that the contours were displaying radio emission and not noise.
After visual inspection of the cutouts of the Bilton et al. (2020) sample of 2,298 galaxies (1670 from non-mergers, 628 from mergers), we find that 191 had a radio detection. These radio detection images were uploaded to Zooniverse222Zooniverse website for visual morphological classification by five team members (M.R., E.M., Y.A.G., L.E.B., J.C.S.P.).
Creating a classification scheme was an iterative process. There was an initial, baseline classification scheme that was created during preliminary work based on the spatial distribution, extension, and strength of the radio emission of the source. After testing this initial classification scheme, we converged on a final morphology scheme that we used on Zooniverse and is as follows (in order of increasing level of disturbance):.
Weak Compact (WC): Emission is (a) roughly circular, (b) on a scale of less than or roughly equal to 12” (2 times the resolution of LoTSS), and (c) has 2 or 3 contours (8-16 detection).
Strong Compact (SC): Emission that is (a) roughly circular, (b) on a scale of less than or roughly equal to 12” (2 times the resolution of LoTSS), and (c) has 4 or more contours ( 32 detection). There may be sources that have 4 contours and are roughly circular but are larger than 12” (up to 30” in size); these are still strong compact 333This condition is meant to catch strong compact sources that only differ from the original classification by being larger than 2 times the LoTSS resolution..
Compact Extended (CE): Emission has a compact (roughly circular and with a size of less than or roughly equal to 30”) core with a singular source of extended emission that is on the order of or greater than 12” (2 times the resolution of LoTSS).
@@
Extended (E): Emission that is (a) clearly on a much larger scale than that of the compact sources on the order of or much greater than 5 times the resolution of LoTSS ( 30”) but is not an FRI, FRII, Bent Tail, or Head Tail source and (b) is irregular in shape (not circular).
@@
Head Tail (HT): An extreme bent tail source where the two jets/tails are indistinguishable
Undetermined444In the Zooniverse project this classification was called “Uncertain” but the name of the classification was changed to Undetermined for the written to avoid confusion with other types of uncertainty. The change in classification class name does not change any results.: There is obviously a source in the anticipated location but the classification is uncertain.
@@
Sources are classified as Undetermined if their morphology does not follow the definition of any provided classifications. The classification system describes the majority of our sample. We note that no morphological classification scheme will be able to exhaustively describe all of the possible complex morphologies. Exemplary sources of each classification are shown in Figure 1.
@@ 3.3 Zooniverse Project
To obtain robust classifications, five team members visually classified each source. For image sharing and collaborative classification, we made use of the Zooniverse Website, a platform, a widely recognized online tool designed for projects reliant on classification-based tasks (Lintott et al., 2008). We established a private project and invited collaborators who are experts in fields relating to the scientific scope of this paper to participate. Upon joining the project, each classifier was presented with a randomly selected radio image from the dataset and prompted to classify it based on the classification criteria described in 3.2. The classifiers were blind to the dynamical state of the parent cluster (merging vs. non-merging) and to which cluster each source came from. The accumulated results from these five evaluators were incorporated into this study for analysis. We note that each classifier was allowed to give each source at most one radio classification and multiple classifications were not allowed.
@@ 3.4 Consensus level
We aggregate the votes from Zooniverse. Following the aggregation of votes from Zooniverse, we define multiple levels of consensus. Consensus level was determined by the number of classifiers (of the 5 total) that agreed on a source classification. For example, in order to create a group of sources for which we were extremely confident in their classification, we define “consensus level 4” as 4 or 5 out of 5 classifiers agreed on a classification. Next, to retain a majority classification but include as many sources as possible, we define “consensus level 3” as 3 of the 5 classifiers agreed on a classification. Consensus level 3 also includes classifications that had a consensus level of 4. Lastly, we define a “non-consensus” class where 2 or fewer classifiers agreed on a classification. There are 71 consensus level 5 sources, 120 consensus level 4 sources, 166 consensus level 3 sources, and 25 non-consensus sources.
We find that the results in this work remain consistent when using consensus level 3 or 4. A comparison between these two consensus levels is described in Appendix B. Because the overall trends are the same in both samples, we use the consensus level of 3 to serve as the standard consensus level for the analysis in this work.
There were 9 sources out of the total 166 consensus level 3 sources that were removed from the sample. There were 6 poor-image-quality sources that were removed. Any sources that appeared to have significant artifact contamination or very low strength as compared to the background image, were classified as bad data (1) or no detection (5) respectively and removed. Additionally, we omit the single source classified as undetermined in the consensus level 3 sample as we want to investigate sources with a distinct radio morphology. We omit these sources (7) from the analysis for a total of 159 consensus level 3 sources. However, for the sake of completeness, we note here that after the radio-AGN delineation process we removed two Extended sources (the reason for which is discussed at length at the end of Section 3.5) for a total of 157 sources. We refer to these sources as the “Zooniverse results” and provide a catalog of these 157 sources with this work for download and display the first 20 rows of this catalog in Table 4.
@@
We further filter the 159 Zooniverse consensus 3 level sources from §3.4 to only those that are radio-AGN in the following way. For radio galaxies with jetted and disturbed morphologies (11 total from the FR I, FR II, FRM, HT, and BT categories) it is clear that the radio emission is the result of an AGN. However, for compact radio sources where the AGN jet is unresolved, we must consider another possibility for the radio emission. In galaxies that do not host an AGN, star-formation produces radio emission, primarily as the result of synchrotron emission from supernovae shocks and thermal emission from H ii regions (Condon, 1992). Generally, radio emission from star-formation is relatively low luminosity, dominating the local radio luminosity function at (Best et al., 2005). In order to analyse the impact of the cluster environment on radio-AGN, it is therefore important to establish that the radio emission for compact objects is indeed the result of an AGN and not star-formation.
@@
For the simple-morphology radio sources in our sample, namely those classified as WC, SC, or CE (141 sources), we estimate the expected MHz emission due to star-formation based on their measured SFRs using the relation of Gürkan et al. (2018). For these sources the MHz flux density is measured from the LoTSS images using Mohan & Raffergy’s PyBDSF (Mohan & Rafferty, 2015). Additionally, the galaxies in our sample are in the SDSS DR8 spectroscopic catalog and have star-formation rate (SFR) measurements from the MPA-JHU catalog (Brinchmann et al., 2004). By comparing the measured radio luminosity, , to the radio luminosity expected from star-formation, , we can thus identify the likely source of the radio emission in our galaxies (see Figure 2). The Gürkan et al. (2018) relation between SFR and radio luminosity has a typical uncertainty of 1.68. To account for this, and given that we are looking to only include radio-AGN in our analysis, we adopt a conservative approach and consider sources with as radio-AGN (red circles on Figure 2). We find that 103 of our simple morphology radio sources have radio emission that can be explained by star formation, leaving us with 38 radio-AGN in our sample that have WC, SC, or CE morphologies.
@@
Further, of the seven E morphologies, three were easily delineated as radio-AGN or star forming using the flux cutoff from their LoTSS catalog value. Two sources are visually determined to possess a degree of jetting on this second pass and therefore are classified as radio-AGN. We honor the original classification of E for these sources, as poorly resolved jetted sources classified as E is a possible uncertainty discussed later in Section 4.1. Of the remaining two E sources, one did not have a LoTSS catalog entry and could not be detected by Mohan & Raffergy’s PyBDSF. Because this source does not a reliable flux measurement and cannot be placed on the Gürkan et al. (2018) relation, we exclude this source. Through optical counterpart analysis, the last E source appears to be the overlap of two spiral galaxies but with only one LoTSS entry. Because it is not possible to determine which galaxy is contributing to the radio emission, we also drop this source from the analysis. Since these two sources were not able to be classified as a radio-AGN or not, we entirely remove them from the sample, reducing the sample from 159 sources to 157 sources.
@@
Using this analysis to separate out the radio-AGN, we find that 43 sources are radio-AGN that have WC, SC, CE or E morphologies in consensus level 3. Thus, as a result, from the 157 Zooniverse sources, the final sample of sources that we use for analysis in the following work is 54 radio-AGN consensus level 3 sources (11 FR I, FR II, FRM, HT, or BT + 43 WC, SC, CE or E).
@@
While ram pressure stripping can boost the radio luminosity of a source, this is most extreme (approximately a factor of 3) at W/Hz, with much lower levels of radio luminosity boosting at W/Hz (see Figure 9 of Roberts et al. 2021). All of our selected radio galaxies have W/Hz and thus subject to low levels of radio luminosity boosting by RPS. Furthermore, our radio galaxy sample are chosen to have radio luminosities of at least 3 times the expected radio luminosity due to SF (see Fig. 2). As such we are confident that all of our selected radio galaxies host a radio-loud AGN.

### v3

@@ 2.2 LoTSS Data
To examine the radio characteristics of the cluster members, we used data from the LOFAR Two-meter Sky Survey (LoTSS: Shimwell et al. 2017). LoTSS provides a resolution of 6′′ with a central frequency of 144 MHz and a median RMS sensitivity in public data of 83 Jy beam-1. We made use of the current public data release of LoTSS which is Data Release 2 (Shimwell et al., 2022); in addition we made use of data processed since that release, which will be made publicly available in LoTSS Data Release 3.
Out of the 33 clusters from Bilton et al. (2020), 13 were covered by the LoTSS survey, 9 nonmerging and 4 merging. Of these 13 clusters, one, A119, was only half covered by the SDSS and was therefore discarded from the analysis. As a result, our final sample consisted of 12 clusters fully covered by LoTSS and SDSS (see Table 1). For each of these clusters we made a mosaicked image that combined all of the available LoTSS pointings covering that cluster, maximizing the sensitivity of the images used for this study.
@@ 2.3 Completeness and Redshift Distribution
It is known that SDSS completeness drops to 65% in denser fields such as a galaxy cluster (Yoon et al., 2008), due to there being more targets than available fibers. To investigate the completeness of our sample, we compare the average mass of the galaxy clusters and the number of cluster members across both dynamical states. When examining the Bilton et al. (2020) SDSS catalogs of the LoTSS covered clusters, we find an average of 62 members per merging cluster and 82 per nonmerging cluster. When we match these averages with average cluster halo mass, we find that mergers have 5.4 x and nonmergers have 6.3 x. Cluster mass, , is determined in Bilton et al. (2020) from the density profile , where is specifically 200 times the critical density of the flat Universe. If we compare the number of cluster members and halo masses of the clusters, we find that merging clusters relative to nonmergers have about 75% of the members and 80% of the mass. The similarity of these values is suggestive that there is not a significant selection bias.
In addition to survey completeness, we investigate the redshift distribution within each merging and nonmerging populations to ensure that they are well matched. Given this work is rooted in morphology, there exists a potential bias of a morphology class being mistaken for another owing to resolution issues coming from redshift differences. For the nonmergers, we find the mean , median , and interquartile range to be 0.06, 0.07, and 0.05, respectively. For mergers, we find the mean , median , and interquartile range to be 0.06, 0.06, and 0.05, respectively. The redshifts are well matched between the cluster dynamical state populations, and we do not expect redshift to bias morphological classifications within each population.
@@ 3 Methods
In order to determine whether or not the merging cluster environment affects the radio morphology of AGN and cluster galaxies, the radio sources need to be classified according to their morphology. To do this, we created images (§3.1), devised a classification system (§3.2), visually classified the radio sources using Zooniverse (§3.3), aggregated the votes from visual inspection and defined consensus levels (§3.4), determined which sources were radio AGNs (§3.5), and performed statistical analyses (§3.6).
@@ 3.1 Cutouts
Since there were a large number of sources and respective cutouts to be made, there was a necessity to be able to create bulk cutouts for the classification process. Thus, we calculated and used a “global” cluster noise (henceforth referred to as RMS or ) for generating contours for the radio sources. To determine the global RMS for each LoTSS image, we made use of SAOImageDS9 (Joye & Mandel 2003). For each mosaicked cluster image, we selected an emission-free region and drew a circle 4′ in diameter within this region. From this circle, we extracted the RMS in mJy beam-1 from the statistics calculated automatically in DS9. For each cluster, we used this global cluster RMS to produce the contours for the cluster member cutouts (reported in Table 1).
The method of calculating the RMS and contour levels displayed for the cutouts can influence the appearance and thus the morphology classification of the source. Given our dataset, which comprises over 1000 sources, we were confronted with a trade-off between efficiency and sensitivity. We investigate whether the morphology classification of a source is RMS (and corresponding contour levels) dependent by selecting a handful of sources (14) showing faint emission potentially related to the source but not captured with contours and testing the effect of using different RMS values. For these sources, we created new cutouts using varying RMS and contour levels. We compared the original, global RMS and contour level for each image to 1) the local RMS (determined from an ′′ circle of an emission-free region in the cutout) with the original 4 2n base for contours; 2) the original global (cluster) RMS but with a 3, 4, 5, 8 contour levels; and 3) the local RMS with 3, 4, 5, 8 contour levels. We find no significant deviation from the original classification results when comparing the new RMS and contour levels. As a result, we are confident that our choice of using a global RMS is robust.
To aid with the classification process, we displayed contours in the cutouts that were log-based starting with a 4 base determined by the cluster RMS, as described above. Thus, the contours begin with 4 and increase by factors of 2n where 1, 2, 3, etc. To optimize visualization for many hundreds of images, we employed a square root scaling technique in aplpy and scaled the images as a percentage of the maximum pixel value (pmax = 100). To see all of these parameters in use, see Figure 1.
@@ 3.2 Radio Morphology Classifications
After the cutouts were made, one team member visually categorized the sources as having a radio detection or nondetection. A source was defined as having a radio detection if it had 8 detection at the cluster member coordinate from the Bilton et al. (2020) catalog, which manifested visually as two contours (at a 4 base). We chose 8 as the threshold for a detection for the following reasoning. Because we used a global, cluster RMS to create the contours for the cluster members, there are a few cases where 4 contour encapsulates the noise instead of the morphology of the source, making the radio morphology of the source difficult to classify. Since we were classifying the sources as detection or nondetection visually (to be consistent with the visual radio morphology classification), we increased the base of detection to the next displayed contour, which in this case is 8 (because we define the contours as 4 2n where 0, 1, 2, etc.) in order to be confident that the contours were displaying radio emission and not noise.
After visual inspection of the cutouts of the Bilton et al. (2020) sample of 2298 galaxies (1670 from nonmergers, 628 from mergers), we find that 191 had a radio detection. These radio detection images were uploaded to Zooniverse222Zooniverse website for visual morphological classification by five team members (M.R., E.M., Y.A.G., L.E.B., J.C.S.P.).
Creating a classification scheme was an iterative process. There was an initial, baseline classification scheme that was created during preliminary work based on the spatial distribution, extension, and strength of the radio emission of the source. After testing this initial classification scheme, we converged on a final morphology scheme that we used on Zooniverse and follows (in order of increasing level of disturbance):
Weak Compact (WC): Emission (a) is roughly circular, (b) is on a scale of less than or roughly equal to 12′′ (2 times the resolution of LoTSS), and (c) has two or three contours (8-16 detection).
Strong Compact (SC): Emission that (a) is roughly circular, (b) is on a scale of less than or roughly equal to 12′′ (2 times the resolution of LoTSS), and (c) has four or more contours ( 32 detection). There may be sources that have 4 contours and are roughly circular but are larger than 12′′ (up to 30′′ in size); these are still strong compact.333This condition is meant to catch strong compact sources that only differ from the original classification by being larger than 2 times the LoTSS resolution.
Compact Extended (CE): Emission has a compact (roughly circular and with a size of less than or roughly equal to 30′′) core with a singular source of extended emission that is on the order of or greater than 12′′ (2 times the resolution of LoTSS).
@@
Extended (E): Emission that (a) is clearly on a much larger scale than that of the compact sources on the order of or much greater than 5 times the resolution of LoTSS ( 30”) but is not an FR I, FR II, bent-tail, or head-tail source and (b) is irregular in shape (not circular).
@@
Head Tail (HT): An extreme bent-tail source where the two jets/tails are indistinguishable.
Undetermined444In the Zooniverse project this classification was called “Uncertain,” but the name of the classification was changed to Undetermined for this paper to avoid confusion with other types of uncertainty. The change in classification class name does not change any results.: There is obviously a source in the anticipated location, but the classification is uncertain.
@@
Sources are classified as “Undetermined” if their morphology does not follow the definition of any provided classifications. The classification system describes the majority of our sample. We note that no morphological classification scheme will be able to exhaustively describe all of the possible complex morphologies. Exemplary sources of each classification are shown in Figure 1.
@@ 3.3 Zooniverse Project
To obtain robust classifications, five team members visually classified each source. For image sharing and collaborative classification, we made use of Zooniverse, a widely recognized online tool designed for projects reliant on classification-based tasks (Lintott et al., 2008). We established a private project and invited collaborators who are experts in fields relating to the scientific scope of this paper to participate. Upon joining the project, each classifier was presented with a randomly selected radio image from the dataset and prompted to classify it based on the classification criteria described in 3.2. The classifiers were blind to the dynamical state of the parent cluster (merging vs. nonmerging) and to which cluster each source came from. The accumulated results from these five evaluators were incorporated into this study for analysis. We note that each classifier was allowed to give each source at most one radio classification and multiple classifications were not allowed.
@@ 3.4 Consensus Level
We aggregate the votes from Zooniverse. Following the aggregation of votes from Zooniverse, we define multiple levels of consensus. Consensus level was determined by the number of classifiers (of the five total) that agreed on a source classification. For example, in order to create a group of sources for which we were extremely confident in their classification, we define “consensus level 4” as four or five out of five classifiers agreeing on a classification. Next, to retain a majority classification but include as many sources as possible, we define “consensus level 3” as three of the five classifiers agreeing on a classification. Consensus level 3 also includes classifications that had a consensus level of 4. Lastly, we define a “nonconsensus” class where 2 or fewer classifiers agreed on a classification. There are 71 consensus level 5 sources, 120 consensus level 4 sources, 166 consensus level 3 sources, and 25 nonconsensus sources.
We find that the results in this work remain consistent when using consensus level 3 or 4. A comparison between these two consensus levels is described in Appendix A. Because the overall trends are the same in both samples, we use the consensus level of 3 to serve as the standard consensus level for the analysis in this work.
There were 9 sources out of the total of 166 consensus level 3 sources that were removed from the sample. There were six poor-image-quality sources that were removed. Any sources that appeared to have significant artifact contamination or very low strength as compared to the background image were classified as bad data (one) or no detection (five), respectively, and removed. Additionally, we omit the single source classified as undetermined in the consensus level 3 sample, as we want to investigate sources with a distinct radio morphology. We omit these sources (seven) from the analysis, for a total of 159 consensus level 3 sources. However, for the sake of completeness, we note here that after the radio AGN delineation process we removed two extended sources (the reason for which is discussed at length at the end of Section 3.5), for a total of 157 sources. We refer to these sources as the “Zooniverse results,” and provide a catalog of these 157 sources with this work for download (see Appendix B).
@@
We further filter the 159 Zooniverse consensus 3 level sources from §3.4 to only those that are radio AGN in the following way. For radio galaxies with jetted and disturbed morphologies (11 total from the FR I, FR II, FRM, HT, and BT categories) it is clear that the radio emission is the result of an AGN. However, for compact radio sources where the AGN jet is unresolved, we must consider another possibility for the radio emission. In galaxies that do not host an AGN, star-formation produces radio emission, primarily as the result of synchrotron emission from supernova shocks and thermal emission from Hii regions (Condon, 1992). Generally, radio emission from star-formation is relatively low luminosity, dominating the local radio luminosity function at (Best et al., 2005). In order to analyze the impact of the cluster environment on radio AGN, it is therefore important to establish that the radio emission for compact objects is indeed the result of an AGN and not star-formation.
@@
For the simple-morphology radio sources in our sample, namely those classified as WC, SC, or CE (141 sources), we estimate the expected MHz emission due to star-formation based on their measured SFRs using the relation of Gürkan et al. (2018). For these sources the MHz flux density is measured from the LoTSS images using Mohan & Raffergy’s PyBDSF (Mohan & Rafferty, 2015). Additionally, the galaxies in our sample are in the SDSS DR8 spectroscopic catalog and have SFR measurements from the MPA-JHU catalog (Brinchmann et al., 2004). By comparing the measured radio luminosity, , with the radio luminosity expected from star-formation, , we can thus identify the likely source of the radio emission in our galaxies (see Figure 2). The Gürkan et al. (2018) relation between SFR and radio luminosity has a typical uncertainty of 1.68. To account for this, and given that we are looking to only include radio AGN in our analysis, we adopt a conservative approach and consider sources with as radio AGN (red circles in Figure 2). We find that 103 of our simple-morphology radio sources have radio emission that can be explained by star formation, leaving us with 38 radio AGN in our sample that have WC, SC, or CE morphologies.
@@
Further, of the seven E morphologies, three were easily delineated as radio AGN or star forming using the flux cutoff from their LoTSS catalog value. Two sources are visually determined to possess a degree of jetting on this second pass and therefore are classified as radio AGN. We honor the original classification of E for these sources, as poorly resolved jetted sources classified as E is a possible uncertainty discussed later in Section 4.1. Of the remaining two E sources, one did not have a LoTSS catalog entry and could not be detected by Mohan & Raffergy’s PyBDSF. Because this source does not have a reliable flux measurement and cannot be placed on the Gürkan et al. (2018) relation, we exclude this source. Through optical counterpart analysis, the last E source appears to be the overlap of two spiral galaxies but with only one LoTSS entry. Because it is not possible to determine which galaxy is contributing to the radio emission, we also drop this source from the analysis. Since these two sources were not able to be classified as a radio AGN or not, we entirely remove them from the sample, reducing the sample from 159 sources to 157 sources.
@@
Using this analysis to separate out the radio AGN, we find that 43 sources are radio AGN that have WC, SC, CE or E morphologies in consensus level 3. Thus, as a result, from the 157 Zooniverse sources, the final sample of sources that we use for analysis in the following work is 54 radio AGN consensus level 3 sources (11 FR I, FR II, FRM, HT, or BT + 43 WC, SC, CE or E).
@@
While ram pressure stripping can boost the radio luminosity of a source, this is most extreme (approximately a factor of 3) at W Hz-1, with much lower levels of radio luminosity boosting at W Hz-1 (see Figure 9 of Roberts et al. 2021). All of our selected radio galaxies have W Hz-1 and thus are subject to low levels of radio luminosity boosting by ram pressure stripping. Furthermore, our radio galaxy sample is chosen to have radio luminosities of at least 3 times the expected radio luminosity owing to star formation (see Fig. 2). As such, we are confident that all of our selected radio galaxies host a radio-loud AGN.